ADS installed miniskirt with retention ring out not calibrated in distance (all out) but clamped and bolted.

Gain ramping slightly improved: EXT:as TS3, INT:start oscillating 0.11.

It was tried to switch off the clean room cooling fans with no changes.

Continuing with FF calibration: removed for no linearity behaviors act DSP#129 and 448

calibrated ff matrix: /mnt/work/idl/adsec672a/meas/ff_matrix/2009_10_20/ff_matrix_+6.005e-05_1.300e-06a_60um.sav

reloaded and calibrated modal: /mnt/work/idl/adsec672a/meas/ff_matrix/2009_10_20/ff_matrix_ff_3.400e-06m_60um_a.sav

removed actuatorDSP# 392

calibrated modal: /mnt/work/idl/adsec672a/meas/ff_matrix/2009_10_20/ff_matrix_ff_3.400e-06m_60um_b.sav

written procedure "vedi_oscillazione.pro" to have a look to the modes that are oscillating: it seems that the modes are in a certain way related to the central hole (coupling between siggnen and ground due to the reduced space between the board on the sheell routing the reference wave and the pipe glued on the reference body?)

-- MarcoXompero - 2009-10-20
Topic revision: r4 - 20 Oct 2009, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback